LUCI1/2 Commissioning

AO-Com

LUCI2 Diffraction-limited Commissioning plan
LUCI2 Diffraction-limited Commissioning plan v2.4 as pdf
Simones outline of LUCI2+FLAO commissioning in January 2015 as pdf
Commissioning Runs in 2015
Email Threads To Save
Observing material
Related material for planning meetings

SL-Com

LUCI2 Seeing-limited Commissioning plan
LUCI2 Commissioning summary logs Nov 1 2013 - Feb 4 2014
Outline LUCI2 closed-dome activities Oct 24 - Nov 15
Available scripts for closed-dome activities LUCI2
Available scripts for seeing-limited night-time commissioning activities LUCI2
Useful stuff for commissioning LUCI2
LUCI2 bugs to be fixed

Recent stuff

Science Night Interrupt Guidelines
Revised LUCI1/2 Diffraction-limited Commissioning plan for 2016B as pdf

Done

Nov 2017 LUCI1/2 AO Commissioning #12

29 November 2017 UT No AO-com due to clouds, non-AO AFC-tests on sky
30 November 2017 UT Closed all night. Arcetri AO & LUCI AO-spec closed-dome tests.
1 December 2017 UT Closed all night. Arcetri AO & LUCI AO-spec closed-dome tests.
2 December 2017 UT Closed all night. LUCI AO-spec closed-dome tests.
3 December 2017 UT First AO AFC-spec tests on sky, half night lost due to weather
9 December 2017 UT AO-tests successfull, more AO AFC-spec tests, seeing-limited AFC-spec tests successfull
10 December 2017 UT Closed all night due to high winds
11 December 2017 UT Clouded out most of the night, a few hours with successfull seeing-limited AFC-spec tests
12 December 2017 UT A tiny bit of PSF-wobble data, some more seeing-limited AFC-spec tests, mostly poor observing conditions

Oct 2017 LUCI1/2 AO Commissioning #11

29 October 2017 UT AO-com AFC-imaging at low elevation completed, AOs worked 5h in parallel perfectly (2nd half), lost 3.5h due to AO errors (1st half)
31 October 2017 UT Closed all night
1 November 2017 UT Clouds at the first few hours, closed due to humidity for the rest of the night
2 November 2017 UT Some AFC-data at medium and high elevation, some unkown LUCI_SW issues, AOs worked relatively reliable
3 November 2017 UT Quite some AFC-data at medium and high elevation, AFC-ima com largely complete (pending analysis)
4 November 2017 UT Closed all night

Sep 2017 LUCI1/2 AO Commissioning #10

19 September 2017 UT AO-tests by AO-teams successfull, 40% of night lost to weather, no LUCI-AO
20 September 2017 UT Bad seeing all night (up to 5" on DIMM), a tiny bit of AO-work, no LUCI-AO
21 September 2017 UT Good progress on AO-side during 1st part of the night, closed 2nd part of night, no LUCI-AO
22 September 2017 UT Most of night lost due to wind, some hours on sky for SA-training and AO-troubleshooting practicing, no LUCI-AO
23 September 2017 UT Closed due to high humidity all night, no LUCI-AO
24 September 2017 UT Lots of AO-tests, LUCI-AO attempt, lots of AO-issues, FLAO(s) very fragile
25 September 2017 UT Bin = 1 / 400 modes seems to work, AFC-ima started successfully, AO-failure rate about 1 / hour

Jun 2017 LUCI1/2 AO Commissioning #10

Cancelled due to Frye fire

Mar 2017 LUCI1/2 AO Commissioning #9_extended

20 March 2017 UT Severe AO@SX-issues, Bino-AO dropped, LUCI2/AO looks promising
21 March 2017 UT AO-issues at SX and DX-side: commissioning stopped

Jan 2017 LUCI1 AO Commissioning #9

14 January 2017 UT Clouded out, LUCI1/2 instrument tests
15 January 2017 UT Clouded out, compensation mirror issues, few AFC-tests
16 January 2017 UT Snow, internet shutdown on Campus all night, no remote obs, AFC-tests from guesthouse
17 January 2017 UT Ice on dome, AFC-tests
18 January 2017 UT Ice on dome, AFC-tests
19 January 2017 UT AO communication issues

November 2016 LUCI1 AO Commissioning - done by LBTO

23 November 2016 UT LUCI1 AO-ima com, FLAO-tests, nice results
24 November 2016 UT LUCI1 AO-ima com, FLAO-tests, nice results

October 2016 LUCI1 AO Commissioning - ARGOS-Com Interrupts (Commissioning #8)

18 October 2016 UT FLAO-SX checkout successfull
19 October 2016 UT First LUCI-1-AO tests, lots of bugs/bugfixing
20 October 2016 UT Workaround for Bayside stage timeouts and large offsets works
21 October 2016 UT More bugs, loop opened 7 times (arbitrator issiues)
22 October 2016 UT Excellent observing conditiions, no progress (three AO issues), fire alarm and other things
23 October 2016 UT 1st useful LUCI-AO com data in 2016B
24 October 2016 UT AFC-tests, several AO-issues
25 October 2016 UT Closed-dome tests
26 October 2016 UT PSF-variability tests, telescope service issues

September 2016 LUCI1/2 Nighttime Diffraction-limited Commissioning #7

Carried out during period September13 - 20 and 22 - 27. Up to 6N are available which
are used depending on the ambient conditions - same guidelines as for the Com-interrupts
earlier this year.

No data taken, 5 good nights were used for FLAO-checkout and troubleshooting, the rest was bad weather

June 2016 LUCI1 AO Commissioning - Science Night Interrupts (Com #6 extention)

14 June 2016 UT LUCI-1-AO tests
UT20160531 - No data taken...problems getting AO system running
UT20160530 - No data taken...clouds rolled in right after interrupt

April 2016 LUCI1 Nighttime Diffraction-limited Commissioning #6

28 Apr 2016 UT Night lost due to bad weather & 2 MOS errors
27 Apr 2016 UT A bit AFC quantification, 2/3 of night for LBTO engineering (pointing model)
26 Apr 2016 UTNight basically lost due to weather and/or adverse observing conditions
25 Apr 2016 UT AFC-tests broad-band filters,. Bino-script phot std, 2/3 lof night lost due to weather
24 Apr 2016 UT Night basically lost due to adverse observing conditions and technical issues
23 Apr 2016 UT AO-tests complete, first successfull ACF-tests on sky,
22 Apr 2016 UT Tests of AO-system all night, seems to be in good shape, optical gain needs a bit of improvement

March 2016 LUCI1 Nighttime Diffraction-limited Commissioning #5

22 Mar 2016 UT 2/3 of night lost due to wind, a bit of AFC and FLAO-testing
21 Mar 2016 UT Successfull night, LUCI1 performed excellently, the AO was not lagging much
20 Mar 2016 UT LUCI1 AO-com has finally started, quite some progress on both sides
19 Mar 2016 UT AO presets from LUCI1 OK, spend all night in collecting data for improving Strehl
18 Mar 2016 UT LUCI1/2 engineering night

May - July 2015 Nighttime Diffraction-limited Commissioning #4/ Seeing-limited checkout Logs

07 July 2015 UT SL testing
01 June 2015 UT AO image quality and PSF-tests, some time lost due to clouds
31 May 2015 UT Closed all night due to ashes in air swirling around LBT
30 May 2015 UT SL-test cont'd, AO-com started
29 May 2015 UT SA training, stability-tests LUCI2 SW cont'd, Fire (humidifier) Alarm
28 May 2015 UT A wide variety of imaging and spectroscopic programs successfully be executed
27 May 2015 UT SL ima- and spec-tests, bugs found, some removed, lots of suggestions

May 2015 Nighttime Diffraction-limited Commissioning #3 Logs

10 May 2015 UT AFC in ima mode works, ZPs N30, TCS/Aladin issues
9 May 2015 UT 3/4 of night lost due to bad weather, TCS issues, some ghost data
8 May 2015 UT 2/3 of night lost due to bad weather, LS-acq. in AO-mode works
7 May 2015 UT Startup AO, AO stage transform, photo ZPs N30 camera
6 May 2015 UT Bad weather all night, camera-repro and ACF-tests
5 May 2015 UT Bad weather all night, repro- and ACF-tests

April/May 2015 Daytime Seeing-limited Commissioning Logs

April 30 - May 5 2015 UT ACF and repro-tests

March/April 2015 Nighttime Diffraction-limited Commissioning #2 Logs

5 April 2015 UT Night mostly lost due to clouds, minor progress
4 April 2015 UT AO successfully operated by LBTO personnel, AO-tests on sky, network issues
3 April 2015 UT LUCI1-tests with new SW, imaging/LS-spectroscopy works, MOS-spectroscopy not
2 April 2015 UT Night was converted to engineering night, some detector/grating tests
1 April 2015 UT Night was converted to engineering night, some detector tests
31 March 2015 UT LUCI2/AGw in lab, Calibration of new gratings
30 March 2015 UT First LUCI AO-script (preset & offsets) executed, first AO-spectrum, power outage
29 March 2015 UT Problem with AGW/AO electronics not solved, SW-Patch made to observe, some AO-tests

March 25-27 2015 Daytime Diffraction-limited commissioning Logs

27 March 2015 UT Problem with AGW/AO electronics
26 March 2015 UT AO malfunction grabber CCD#47 resolved
25 March 2015 UT Pupil alignment, image scale N30 camera, AO malfunction grabber CCD#47

January 2015 Nighttime Diffraction-limited Commissioning #1 Logs

27 January 2015 UT Closed all night, various NCPA and vibration tests
28 January 2015 UT Closed all night, NCPA Rotator angle now fully covered
29 January 2015 UT 2hrs useful, first AO-images with LUCI+FLAO on sky, strehl up to 75%
30 January 2015 UT Closed all night, NCPA tests, check N30 masks
31 January 2015 UT Closed all night, NCPA tests (Strehl up to 80% now), SW tests
01 February 2015 UT Closed all night

January 19-25 2015 Daytime Diffraction-limited commissioning Logs

19 January 2015 UT Tasks #0, #1 completed, Tasks #4 and 8 partly done
20 January 2015 UT ND-mount installation successfull, MOS error but solved
21 January 2015 UT Work on NCPA (Task#3), image scale N30-camera, GEIRS bug
22 January 2015 UT Strehl lower than expected, GEIRS windowed readout works
23 January 2015 UT Work on NCPA (Task#3), ND-mount removed
24 January 2015 UT Work on NCPA (Task#3), Strehl in H-band > 75% now
25 January 2015 UT Even more work on NCPA (Task#3)

February 2014 Daytime Seeing-limited Commissioning Logs

1-28 February 2014 UT Various tests

January 2014 Nighttime Seeing-Limited Commissioning #3 Logs

22 January 2014 UT 3/4N lost to weather, detector tests, some MOS-tests
21 January 2014 UT Pointing model, LS/MOS-spectroscopy
20 January 2014 UT SW stable, training MOS-acq, deep imaging
19 January 2014 UT Simulating scientific observations (deep imaging/spectroscopy)
18 January 2014 UT Collimation issues, CofR, deep spectra, throughput
17 January 2014 UT AGw-transfrom works, MOS-acq. perfect, lots of functionality and characterisation data
16 January 2014 UT AGw-transform, ZPs, throughput, persistence/crosstalk, LS-acq. tests

December 2013 Nighttime Seeing-limited Commissioning #2 Logs

19 December 2013 UT Offset-tests, throughput gratings, phot ZPs, 7h lost to weather
18 December 2013 UT MOS-acquisition works, AdSec checkout, AGw-verification, IQ/scale/distortion data
17 December 2013 UT AGw-transform data, MOS-acquisition tests
16 December 2013 UT Longslit acquisition works, MOS-issues
15 December 2013 UT AGw-transform data, LS-acquisition tests, ghosts/reflections, phot ZP's
14 December 2013 UT LS-acquisition tests, 1/3 night lost to humidity
13 December 2013 UT Bad weather, closed-dome tests

November 20 - December 12 Daytime Seeing-limited Commissioning Logs

20 November - 12 December 2013 UT Various tests

November 2013 Nighttime Seeing-Limited Commissioning #1 Logs

15 November 2013 UT LUCI1+2 binocular preset+offset, LS- and MOS-acq. tests, 3hrs lost to weather
14 November 2013 UT Technical issues, bino-preset, first spectra, stability guiding, large offsets
13 November 2013 UT Closed all night, flexure calibration
12 November 2013 UT Photometric ZPs, AGw transformation verification (failed)
11 November 2013 UT AGw transform, telescope focus
10 November 2013 UT 1st night successfull!

October 24 - November 15 2013 Daytime Seeing-limited Commissioning Logs

15 November 2013 UT LUCI2 auxilliary cryostat exchange
7/8 November 2013 UT Pupil alignment & field centering, LUCI1+2 binocular preset, ARGOS imaging & spectroscopy
5 November 2013 UT Flexure data
4 November 2013 UT Flexure data
3 November 2013 UT Some flexure tests
2 November 2013 UT No test done
1 November 2013 UT No tests done
31 October 2013 UT No tests done
30 October 2013 UT Basic tests completed
29 October 2013 UT No tests done
28 October 2013 UT No tests done
25 - 27 October 2013 UT Issue with Rack Cooling Control solved, flexure data scripts work
24 October 2013 UT Basic function tests, MOS tests, issue with the Rack Cooling Control

-- %USERSIG{JochenHeidt - 2017-09-15}%
Topic attachments
I Attachment Action Size Date Who Comment
AOcommPlan19-30jan14V3.pdfpdf AOcommPlan19-30jan14V3.pdf manage 206 K 22 Jan 2015 - 10:04 JochenHeidt Simones outline of the LUCI2+FLAO commissioning January 2015
ComPlan_2016B.pdfpdf ComPlan_2016B.pdf manage 52 K 05 Sep 2016 - 11:51 JochenHeidt LUCI1/2 AO-commssioning plan for 2016B
LBT-LUCIFER-PLA-004.pdfpdf LBT-LUCIFER-PLA-004.pdf manage 205 K 22 Jan 2015 - 10:00 JochenHeidt Commissioning plan LUCI2 v2.4
Topic revision: r208 - 12 Dec 2017, JochenHeidt
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback